The price for a good has risen from 675 rub. to 689 rub. The supply has increased from 3100 units to 3450 units. Calculate the price elasticity of supply? 


1
Expert's answer
2021-09-22T07:24:33-0400

Solution:

Price elasticity of supply (PES) = "=\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}"


PES = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


Q1 = 3,100                                                P1 = 675

Q2 = 3,450                                                P2 = 689


PES = "\frac{3,450 -3,100}{(3,450+3,100)/2 }\div \frac{689 -675}{(689+675)/2 }"


PES = "\frac{350}{3,275} \div\frac{14}{682} = \frac{0.107}{0.021} = 5.21"


Price elasticity of supply (PES) = 5.21

This means that the good has a relatively elastic supply, indicating high responsiveness to changes in price.